# Limits of open ASEP stationary measures near a boundary

### v1

@@ Appendix B Total variation bounds of Askey–Wilson signed measures
We return to the proof. By the reasoning below Claim B.4, we conclude the proof of bound B.2:
$\left\lVert P_{s,t}(x,\mathsf{d}y)\right\rVert_{TV}\leq\frac{K}{(s-t)^{2}}.$
Our last goal is to show that the above constant $K$ can be chosen as not depending on $D$. Since $D\in(-1,0]$, $\max(|a|,|b|,|c|,|d|)$ can be bounded by a constant not depending on $D$. Therefore the total number of atoms $\operatorname{card}\left(F(a,b,c,d)\right)$ as well as the constant prefactors appearing in the atom mass bounds (B.5) can be bounded by constants independent of $D$. In the proof of Claim B.4 (which bound all the linear factors $|1-q^{l}\mathfrak{f}/\mathfrak{e}|$), $D$ does not play a role at all. Therefore the uniform constants chosen from Claim B.4 does not depend on $D$. In summary, constant $K$ can be chosen as independent of $D$. We conclude the proof. ∎
@@ Appendix C Time-reversal of Askey–Wilson signed measures
We prove a special symmetry of multi-dimensional Askey–Wilson signed measures known as the time-reversal. In the case that this symmetry only involves actual probability measures (i.e., on the fan region $AC<1$), this symmetry appears as [4, equation (5.10)]. However the proof is not explained clearly enough therein. Here we adopt a different approach to prove the result for the general case.
@@
Assume $q\in[0,1)$ and $A,B,C,D\in\mathbb{R}$. We have:
@@
Before we begin the proof of Proposition C.1, we recall the orthogonality and the projection formula of Askey–Wilson polynomials and signed measures. The Askey–Wilson polynomials $w_{j}(x):=w_{j}(x;a,b,c,d)$, $j\in\mathbb{N}_{0}$ are defined by three term recurrence in [18] (see also [10, Section 2.1]).
@@
$\pi_{t}^{(A,B,C,D)}(\mathsf{d}x)=\nu\left(\mathsf{d}x;A\sqrt{t},B\sqrt{t},\frac{C}{\sqrt{t}},\frac{D}{\sqrt{t}}\right)=\nu\left(\mathsf{d}x;\frac{C}{\sqrt{t}},\frac{D}{\sqrt{t}},A\sqrt{t},B\sqrt{t}\right)=\pi_{1/t}^{(C,D,A,B)}(\mathsf{d}x).$
@@
In view of (the uniqueness part of) the Riesz representation theorem (see for example [19, Theorem 6.19]), we have

### v2

@@ Appendix B Total variation bounds of Askey–Wilson signed measures
We return to the proof. By the reasoning below Claim B.4, we conclude the proof of (B.2). It is clear from our choice of $\varepsilon>0$ that it is independent of $D$. Our last goal is to show that the constant $K$ can also be chosen as not depending on $D$. Since $D\in(-1,0]$, $\max(|a|,|b|,|c|,|d|)$ can be bounded by a constant not depending on $D$. Therefore the total number of atoms $\operatorname{card}\left(F(a,b,c,d)\right)$ as well as the constant prefactors appearing in the atom mass bounds (B.5) can be bounded by constants independent of $D$. In the proof of Claim B.4 (which bound all the linear factors $|1-q^{l}\mathfrak{f}/\mathfrak{e}|$), $D$ does not play a role at all. Therefore the uniform constants chosen from Claim B.4 does not depend on $D$. In summary, constant $K$ can be chosen as independent of $D$. We conclude the proof. ∎
@@ Appendix C Time-reversal of Askey–Wilson signed measures
We prove a special symmetry of multi-dimensional Askey–Wilson signed measures known as the time-reversal. In the case that this symmetry only involves actual probability measures (i.e., on the fan region $AC<1$), this symmetry appears as [9, equation (5.10)]. However the proof is not explained clearly enough therein. Here we adopt a different approach to prove the result for the general case.
@@
Assume $q\in[0,1)$ and $A,B,C,D\in\mathbb{R}$ satisfying $q^{l}ABCD\neq 1$ for all $l\in\mathbb{N}_{0}$. We have:
@@
Before we begin the proof of Proposition C.1, we recall the orthogonality and the projection formula of Askey–Wilson polynomials and signed measures. The Askey–Wilson polynomials $w_{j}(x):=w_{j}(x;a,b,c,d)$, $j\in\mathbb{N}_{0}$ are defined by three term recurrence in [1] (see also [30, Section 2.1]).
@@
$\pi_{t}^{(A,B,C,D)}(\mathsf{d}x)=\nu\left(\mathsf{d}x;A\sqrt{t},B\sqrt{t},\frac{C}{\sqrt{t}},\frac{D}{\sqrt{t}}\right)\\ =\nu\left(\mathsf{d}x;\frac{C}{\sqrt{t}},\frac{D}{\sqrt{t}},A\sqrt{t},B\sqrt{t}\right)=\pi_{1/t}^{(C,D,A,B)}(\mathsf{d}x).$
@@
In view of (the uniqueness part of) the Riesz representation theorem (see for example [26, Theorem 6.19]), we have